OpenStudy (anonymous):

What is the vertex of this equation f(x) = −x2 + 48x − 432

OpenStudy (triciaal):

vertex form of equation a(x-h)^2 + k h = -b/(2 a ) h = x coordinate of vertex you can use completing the square